OpenStudy (morganchavez):

ALGEBRA 2, UNIT 3, LESSON 5: Graphing Quadratics ALTERNATE ASSIGNMENT Did you know parabolas are all around you in the real world!? If you’ve traveled to the Midwest, you may have seen this one before: The Gateway Arch in St. Louis, Missouri is a 630-foot high monument that was first designed in 1947. Construction of the stainless steel arch lasted many years, and the arch was not completed until October 1965. It was created as a monument to the westward expansion of the United States, and to this day it is the tallest man-made monument in America, as well as the world’s tallest arch

OpenStudy (morganchavez):

Use the following function to complete #1-3 below f(x)=2(x-6)^2+4 1) Using complete sentences, explain how to convert f(x) into standard form. Give the final standard form equation in your answer. 2) Graph f(x) by hand or with technology. Include a photo or screenshot of the parabola. 3) Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (morganchavez):

please help will medal and fan i beg of yall lol

OpenStudy (anonymous):

Hun I honestly would help but I took this stuff back in 9th grade I can't even remember any of it

OpenStudy (morganchavez):

thx :( lol 11 grade math is hard

OpenStudy (morganchavez):

@One098 @onlinestudentxP @gorv @goldendragon01 @Ilovecake @jrc5101520 @jose12345 @Haichi @Jadeyn @StudyGurl14 @Omg_Bmth @nincompoop @nikki__kitty @ganeshie8

OpenStudy (morganchavez):

@Crissy15 @perl @paki @picachu341 @prowrestler @usedapples @Loser66 @lolpizza. @damien @Donblue22 @sammie2017 @dan815 @rockstar0765

OpenStudy (morganchavez):

GUYS PLEASE HELP ME I AM LIKE LITERALLY BEGGING YALL

OpenStudy (morganchavez):

lik i know the standard form of the porabola is y=2x^2-27x+76 which is #1

OpenStudy (morganchavez):

so i mean i can explain tht but idk how to do #3

OpenStudy (morganchavez):

@mathmath333 @magepker728

OpenStudy (paki):

do you know the slope of the line...?

OpenStudy (paki):

@morganchavez

OpenStudy (morganchavez):

no but i can try n find it :)

OpenStudy (paki):

ok, average rate of change will be average rate of change = [f(7) - f(4)] / (7 - 4) ...

OpenStudy (morganchavez):

?

OpenStudy (morganchavez):

um ok so lik vertex means slop right ?

OpenStudy (paki):

yup

OpenStudy (morganchavez):

ok so the slope is (6,4)

OpenStudy (morganchavez):

right ?

OpenStudy (paki):

agree

OpenStudy (morganchavez):

like y do i need the slope ? like what number?

OpenStudy (paki):

slope is equal to the average rate of change...

OpenStudy (morganchavez):

like im so sorry paki but im so confused n i dont know y

OpenStudy (morganchavez):

ok so how would i write it out then ?

OpenStudy (morganchavez):

im so sorry im a visual learner im not trying to make u do my work i swear i jus need help my teacher on flvs doesnt explain nothing n ive been on this for 4 days now

OpenStudy (paki):

rite... so you need all these basics from the start i think.... @morganchavez

OpenStudy (morganchavez):

yes please if u dont mind please

OpenStudy (morganchavez):

well lik not the very beginging jus the part where the average rate part starts

OpenStudy (paki):

where you have the problem...?

OpenStudy (morganchavez):

lik #3 like where n how i put tht equation togehter with the slop n stuff

OpenStudy (paki):

ok... please have a look here... http://openstudy.com/study#/updates/52c35405e4b0fef2cb9fd44e

OpenStudy (morganchavez):

ok thx

OpenStudy (paki):

pleasure...

OpenStudy (morganchavez):

y1=2(7-9)2+17 y2=2(4-9)2+17 right ?

OpenStudy (morganchavez):

@paki so now i jus solve them and by solving i mean simlify right ?